• Martin, Inc., is considering the development of a subsidiary in Sydney, Australia that would manufacture and sell high quality guitars on the local Australian Market. • As one of Martin’s financial managers, you have asked the finance, marketing and manufacturing departments to provide you with all of the relevant input so you can perform a capital budgeting analysis to determine whether to undertake this project or not. • In addition, a contingent of Martin executives have met with Australian government officials in Sydney to discuss the proposed subsidiary. • The project would end in 5 years. All of the relevant information that you need to perform this capital budgeting analysis follows here. • Initial investment: A$12,500,000 million (A$ = Australian dollars) • Price and consumer demand: Year 1 and 2: 20,000 units @ A$650/unit Year 3: 30,000 units @ A$760/unit Year 4: 40,000 units @ A$850/unit Year 5: 40,000 units @ A$880/unit • Costs Variable costs: Years 1 & 2 A$370/unit, Year 3 A$445/unit, Years 4 and 5 A$600/unit Fixed costs: A$2,000,000 per year ( A$1,000,000 lease and A$1,000,000 other fixed expense) • Depreciation A$1,000,000 per year • Tax laws: 30% income tax • Remitted funds: 15% withholding tax on remitted funds • Exchange rates: Spot exchange rate of $US 0.65 for Australian dollar • Salvage values: A$6 million • Required rate of return: 18% Given all of the preceding information, please create an Excel spread sheet that shows the year by year amounts that you have taken into consideration in your capital budget analysis and decide whether or not you will undertake this project.
Answer #1

Whether to undertake the project: the decision will be based on net present value of cash flows where

Net Present value = Present value of cash inflows - present value of cash outflows

Present value of cash inflows = Year-on-year cash flows discounted at required rate of return

Present value of cash outflows = initial cash outlay

Free cash flow calculations

1 2 3 4 5
Price/unit 650 650 760 850 880
Less: Variable cost per unit 370 370 445 600 600
Contribution per unit 280 280 315 250 280
Demand (units) 20000 20000 30000 40000 40000
Contribution 5600000 5600000 9450000 10000000 11200000
less : fixed cost 2000000 2000000 2000000 2000000 2000000
less : depreciation 1000000 1000000 1000000 1000000 1000000
Profit before tax 2600000 2600000 6450000 7000000 8200000
Less: tax @30% 780000 780000 1935000 2100000 2460000
Profit after tax 1820000 1820000 4515000 4900000 5740000
Add: depreciation 1000000 1000000 1000000 1000000 1000000
Free Cash Flow 2820000 2820000 5515000 5900000 6740000
Add: salvage value 6000000
Amount available to remitt 2820000 2820000 5515000 5900000 12740000
Less: Withholding tax on remittance @ 15% 423000 423000 827250 885000 1911000
Amount to be remitted 2397000 2397000 4687750 5015000 10829000
Disounting factor @ 18%                  0.85                  0.72                  0.61                  0.52                  0.44
Discounted cash flow 2031356 1721488 2853109 2586681 4733456
Present value of cash inflows A$ 13926090.30
Exchange rate 0.65
Present value of cash inflows USD 9051958.69

Present value of initial cash outflow = A$ 12500000

Present value of initial cash outflow = 12500000 * .65 = 8125000

Net present value = Present value of cash inflows - present value of cash outflows

= 9051958.69 - 8125000 = 926958.69

Because present value at required rate of return of 18% is positive you will undertake this project.

Note 1 - Excel Calculations

В C E F 4. 1 2 5 2 Price/unit 650 760 650 850 880 Less: Variable cost per unit 3 370 370 445 600 600 4 Contribution per unit

Note 2 - Excel formulas

В C E F 4 1 2. 850 2 Price/unit: 650 650 760 880 Less: Variable cost per unit 370 370 600 445 600 4 Contribution per unit 5 D

Note 3 Assumption

Asset Cost = 12500000

Total depreciation in 5 years = 5000000

Salvage value = 6000000

Value unaccounted for = 12500000-5000000-6000000 = 1500000

We have assumed that this 1500000 has no tax implications in the 5th year